# Data-Driven Prescriptive Analytics Applications: A Comprehensive Survey

### v1

@@ Abstract
Prescriptive Analytics (PSA), an emerging business analytics field suggesting concrete options for solving business problems, has seen an increasing amount of interest after more than a decade of multidisciplinary research. This paper is a comprehensive survey of existing applications within PSA in terms of their use cases, methodologies, and possible future research directions. To ensure a manageable scope, we focus on PSA applications that develop data-driven, automatic workflows, i.e. Data-Driven PSA (DPSA). Following a systematic methodology, we identify and include 104 papers in our survey. As our key contributions, we derive a number of novel conceptual models: In terms of use cases, we derive 10 application domains for DPSA, from Healthcare to Manufacturing, and subsumed problem types within each. In terms of individual method usage, we derive 5 method types and map them to a comprehensive taxonomy of method usage within DPSA applications, covering mathematical optimization, data mining and machine learning, probabilistic modelling, domain expertise, as well as simulations. As for combined method usage, we provide a statistical overview of how different method usage combinations are distributed and derive 2 generic workflow patterns along with subsumed workflow patterns, combining methods by either sequential or simultaneous relationships. Finally, we derive 4 possible research directions based on frequently recurring issues among surveyed papers, suggesting new frontiers in terms of methods, tools, and use cases.
Index Terms:
prescriptive analytics, survey, application, optimization, machine learning, big data.
I Introduction
@@
One of the current frontiers within business analytics is Prescriptive Analytics (PSA), which arrived as a term in the 2010s [5, 6]. Going beyond understanding the past or predicting the future, the purpose of PSA is to suggest concrete courses of actions in the present [7]. With PSA, proactive decision making is supported by combining, e.g., mathematical optimization with data-driven machine learning (ML) [8].
While still being a relatively new field steadily gaining foothold in industry, PSA has seen an increasing level of interest among researchers in recent years [9], with the planned 2025 ACM SIGMOD conference, the most important scientific conference within data management systems, now having PSA as a dedicated topic in their call for papers for the first time ever [10].
With more than a decade of multidisciplinary research activity to comprehend, along with an increasing level of interest among researchers, there is currently a need for synthesis and guidance on how to progress PSA further as a field. Yet as our analysis in Section II shows, no comprehensive survey on PSA applications, providing complete coverage of problem and solution types encountered across all application domains, exists as of yet. This crudely shaped research gap is arguably a curious predicament. As others have previously observed, applications indeed make up the vast majority of the PSA research field in terms of the number of papers published, compared to the number of papers discussing general theoretical or technical issues [11].
The primary objective of this survey is to provide a comprehensive analysis of where the many applications of PSA have been, and where they are going. To ensure a more well-defined, manageable survey scope, we focus on a prominent subset of PSA that we call Data-Driven PSA (DPSA, pronounced ”deepsa”). The pursuit of our primary objective is decomposed into the following three research questions:
  1. RQ1:
@@ 3. RQ3:
  • We derive 10 application domains and subsumed problem types to form a comprehensive, updated overview of current DPSA use cases.
@@
  • We derive 5 method types and a taxonomy of subsumed methods, forming a comprehensive, updated overview of all methods found within existing DPSA applications.
@@
  • We derive 2 generic workflow patterns and subsumed concrete workflow patterns, and provide a statistical overview of the prevalence of different method type usage combinations, forming a comprehensive, updated overview of how DPSA methodologies are structured in current applications.
@@
  • Finally, we derive 4 possible research directions reflecting the current state of the art, based on an analysis of frequently recurring issues within the DPSA application research field.
@@
The remainder of this paper is structured as follows: First, we provide a systematic analysis of how the present survey marks a relevant scientific contribution compared to previous surveys in Section II. Section III covers relevant background for readers unfamiliar with PSA. Then Section IV provides a definition of DPSA, to delimit the scope of this survey. Sections V-VII move on to present the chosen survey methodology, choices made as part of the survey process, along with a statistical overview of included papers, respectively. Section VIII, Sections IX-XI, and Section XII address each research question in turn, providing a comprehensive framework of problem and solution types found within existing application research, along with a number of possible future research directions. We provide concluding remarks in Section XIII.
II Related work
@@
Several surveys about PSA have been published within recent years, as summarized recently by Wissuchek and Zschech [12]. An analysis of feature overlaps between existing surveys and this one can be found in Table I, with papers sorted by publication year from left to right.
Feature / Paper [13] [11] [9] [14] [15] [16] [17] [18] [19] [20] This study
Papers from this decade included
PSA applications the primary focus
All application domains within scope
Analysis of existing application domains (✓) (✓) (✓)
Analysis of problem types per application domain (✓) (✓) (✓)
Analysis of specific PSA methods in use (✓)
Analysis of PSA method workflow patterns
@@
TABLE I: This study’s overlaps with existing PSA surveys, with papers sorted by publication year from left to right. (✓) denotes significantly limited fulfilment. [20] and [21] are respectively older versions of [12] and [9], and are therefore left out for the sake of brevity.
@@
First, we note that a number of existing surveys don’t cover any papers from this decade [13, 11, 9, 14]. As our analysis in Section VII shows, 76.9% of our surveyed papers were published within the 2020-2024 year range, with the number of papers having accelerated significantly over the past three years. Looking solely at the recency of content, this survey therefore provides substantially updated knowledge.
@@
Furthermore, only a few past surveys have had a primary focus on PSA applications, with others focusing on systems and tools [11], methodologies [9, 17], or sociotechnical system analysis [12, 13]. Some surveys indeed focus primarily on applications, but they either limit their scope to studying specific application domains or analysis methods utilized therein [15, 16, 18, 17].
@@
Remaining application-oriented papers limit their scope to a significantly limited subset of applications by more or less explicit survey-methodological choices [14, 19]. One of these papers survey ”various” PSA applications, as its title aptly conveys, yet only 21 in total, and without specifying how these papers were found or selected [14]. The other paper explicitly limits its search query to only include PSA papers intersecting with keywords under the topic of ”sustainable operations research” [19]. These methodological choices, excluding the larger portion of existing PSA papers, severely limit how comprehensive analyses of PSA application domains and problem types can be, which is reflected by parenthesized check marks in Table I. Note that the present survey mainly limits itself to DPSA for definitional reasons - to avoid vague definitions of the PSA field that would conflate it with many adjacent research areas (cf. Section IV). The coverage of this survey is therefore deemed comprehensive with respect to the definition of PSA being followed.
@@
Similarly to the two papers discussed previously, the most recent PSA survey paper [20] only covers the ”more prevalent” (in terms of the number of papers) application domains and problem types, or ”affordances” in the chosen nomenclature, which is also reflected in Table I. The aforementioned survey also mentions a number of overall PSA method categories and example methods, although these are not derived from its own set of survey papers, but from Lepenioti et al.’s survey from 2020 [9].
@@
In addition to being comprehensive unlike previous endeavours, this survey marks a relevant research contribution by introducing novel analysis dimensions, namely an analysis and taxonomy of how solution methods are combined into PSA workflows. Although others have considered the relevance of studying this topic [12], only individual workflow patterns have been discussed fragmentarily within PSA research so far (e.g., [22]).
III Background on PSA
@@
As described in in Davenport and Harris’ seminal book on the subject, Competing on Analytics: The New Science of Winning [23], problem solving being evidence-based more concretely entails that business decisions should be driven by the ”extensive use of data, statistical and quantitative analysis, explanatory and predictive models, and fact-based management”, enabled by a set of technologies and processes [23,  p. 7].
@@
IV PSA definitions and survey scope
To ensure a well-defined scope, it was ultimately deemed necessary to use a more specific version of Definition 2 in this survey. Throughout history, many prescriptive approaches have been used within human organizational endeavours: from ancient astronomy using the constellations to time when to harvest the crops [31,  p. 4], to modern operations research arriving to support military efforts during the Second World War [30]. Even disregarding definitional issues, a single survey paper cannot realistically aspire to cover everything written within the scope of Definition 2 since the dawn of history, let alone entire fields like operations research, which came to be more than 80 years ago [32], with a corresponding number of papers to boot.
Instead, we have chosen to refine our scope to focus on prescriptive approaches following a data-driven, automated paradigm in terms of solution methods. With this focus, we answer recent calls for more data-driven PSA research [10, 9, 33]. As we will argue next, the aforementioned PSA paradigm is furthermore characteristic of the business analytics and technological context in which the ”PSA” term originally arose during the past decade (cf. Section III).
To start, consider the defining characteristics of analytics given in Davenport and Harris’ seminal book [23]. First, analytics is literally understood as a subset of business intelligence: ”a set of technologies and processes that use data to understand and analyse business performance” (our emphasis) [23,  p. 8]. Analytics is thus defined as a fundamentally data-driven endeavour. Second, while it is recognized in the same book that analytics may technically either be input to human decisions, and thus subjected to manual intervention, or be fully automated, it is also stated that any ”rational individual” (our paraphrase) using analytics in the current age would prefer to rely on information technology over pen and paper [23,  p. 8], signifying a preference for automation when possible. Apart from this singular perspective on the matter, many existing definitions of and rationales behind business analytics explicitly mention some combination of being data-driven and leveraging automation as part of its key traits, as summarized in [3].
In addition, note how the ”PSA” term arose within the technological context of the 2010s (cf. Section III): a time period during which increasingly scalable computing systems for (Big) data-driven, automated analytics had a momentous impact on industry and society [34, 35], and a time period during which the data-driven analytics ecosystem underwent rapid expansion in terms of user base and available tools [36], offering increasing levels of workflow automation [37].
@@ IV PSA definitions and survey scope
Data-driven prediction, exemplified by data mining and ML methods, accords with original characterizations of business analytics discussed in this section, while automatic prescription, exemplified by mathematical optimization, accords with how the value proposition of PSA was originally formulated by IBM (cf. Section III): finding the best decision options within problem spaces otherwise too large or complex to search manually.
@@
V Survey methodology
This section outlines the utilized survey methodology, taking its onset in the recommendations followed by previous surveys within PSA [12, 9]. Section VI and VII provides a statistical summary of the search process and identified papers, respectively. Throughout, we follow the PRISMA 2020 guidelines for systematic literature review reporting [38].
V-A Databases
@@ V-B Search queries
V-C Selection criteria
To narrow down initial search results to a subset adequately addressing the research questions (cf. Section I), a set of selection criteria was developed, used for screening the initial set of records in several phases. Within the context of this survey, we will refer to the remaining set of records after screening with these criteria as the set of surveyed papers, or just papers for short when unambiguous. The set of selection criteria is as follows, to be explained in the remainder of this section:
  1. C1:
@@ 6. C6:
As for C5, the requirement of developing an end-to-end application means that the paper must analyze a concrete problem solved with concrete methods. For instance, it is insufficient for a paper to simply discuss hypothetical applications of PSA within some application domain on a purely conceptual level, or to propose ”machine learning” as a generic solution to a specific real-world problem without further details. This criterion is introduced in order to be able to assess adherence to Definition 3 and adequately address the research questions (cf. Section I).
@@
VI Survey process
This section describes the process of how the survey procedure was executed, including choices made during screening. Using the databases, queries and selection criteria of Section V-A, V-B, and V-C, the screening process for obtaining the final set of surveyed papers proceeded as depicted in Figure 2. This figure depicts a flow diagram according to the PRISMA 2020 guidelines for systematic reviews [38], summarizing the number of records and choices made during each screening phase.
Figure 2: Flow diagram of the executed survey procedure, using the PRISMA 2020 format [38]. Each box in the left chain represents a phase in a sequential screening workflow, along with its number of input records. The boxes to the right summarize how many records were excluded in each screening phase, as well as the reasons why, with respect to the selection criteria of Section V-C.
@@ VI Survey process
A few notes on choices made during each phase are in order, and the reader is advised to confer Figure 2 as needed. For the outlet screening phase, records excluded due to C2 were PhD theses, workshop papers, and papers from the Computing Research Repository (CoRR) which were not peer-reviewed. Exclusions due to C3 and C4 followed the objective metrics described in Section V-C by looking up the relevant information in the chosen online registers, with the larger part of records being excluded due to the outlet being low-impact (C4).
@@
VII Survey statistics
A total of 104 papers were obtained from executing the survey procedure. Figure 3 summarizes the distribution of papers per year, which spans the entire 2015-2024 range. There is evidently a pronounced skew towards more recent years, which might be attributed to the increasing popularity and impact of DPSA applications within PSA research, with the 2020-2024 range making up about 76.9% of all papers. Supplementary survey statistics can be found in Appendix A in the electronic appendix.

### v2

@@ Abstract
Prescriptive Analytics (PSA), an emerging business analytics field suggesting concrete options for solving business problems, has seen an increasing amount of interest after more than a decade of multidisciplinary research. This paper is a comprehensive survey of existing applications within PSA in terms of their use cases, methodologies, and possible future research directions. To ensure a manageable scope, we focus on PSA applications that develop data-driven, automatic workflows, i.e., Data-Driven PSA (DPSA). Following a systematic methodology, we identify and include 104 papers in our survey. As our key contributions, we derive a number of novel taxonomies of the field and use them to analyse the field’s temporal development. In terms of use cases, we derive 10 application domains for DPSA, from Healthcare to Manufacturing, and subsumed problem types within each. In terms of individual method usage, we derive 5 method types and map them to a comprehensive taxonomy of method usage within DPSA applications, covering mathematical optimization, data mining and machine learning, probabilistic modelling, domain expertise, as well as simulations. As for combined method usage, we provide a statistical overview of how different method usage combinations are distributed and derive 2 generic workflow patterns along with subsumed workflow patterns, combining methods by either sequential or simultaneous relationships. Finally, we derive 4 possible research directions based on frequently recurring issues among surveyed papers, suggesting new frontiers in terms of methods, tools, and use cases.
keywords:
prescriptive analytics , survey , application , optimization , machine learning , big data
journal: Information Systems
[…]
[1]organization=Department of Computer Science, Aalborg University, addressline=Selma Lagerløfs Vej 300, city=Aalborg Ø, postcode=DK-9220, country=Denmark
1 Introduction
@@
One of the current frontiers within business analytics is Prescriptive Analytics (PSA), which arrived as a term in the 2010s [5, 6]. Going beyond understanding the past or predicting the future, the purpose of PSA is to suggest concrete courses of actions in the present [7]. With PSA, proactive decision making is supported by combining, e.g., most archetypically, mathematical optimization with data-driven machine learning (ML) [8].
Such workflows are especially useful within enterprises where success hinges on a harmonious marriage between prediction and planning. Such is the case in many seemingly unrelated business domains, from advertising to healthcare wherein, e.g., customer and patient treatment plans alike must be weighted by prognostic data [9]. From conversations with our own industrial contacts and collaborators within retail, energy trading, and beyond, we indeed find that many real-world companies are already doing PSA, even though they might be unaware of the ”PSA” term within their organization. To them, it goes without saying that one must of course, e.g., estimate future user demand in order to manage inventory logistics, or likewise, to make bids on the European electricity market a day before delivery.
PSA also has seen an increasing level of interest among researchers in recent years [9], especially due to the inherent technical challenges of combining, e.g., data-driven ML and mathematical optimization technologies [10]. Indeed, the 2025 ACM SIGMOD conference, the most important scientific conference within data management systems, now has PSA as a dedicated topic in their call for papers for the first time ever [11]. With more than a decade of multidisciplinary research activity to comprehend, along with an increasing level of interest among researchers, there is currently a need for synthesis and guidance on how to progress PSA further as a field. Yet as our analysis in Section 2 shows, no comprehensive survey on PSA applications, providing complete coverage of problem and solution types encountered across all application domains, exists as of yet. This crudely shaped research gap is arguably a curious predicament. As others have previously observed, applications indeed make up the vast majority of the PSA research field in terms of the number of papers published, compared to the number of papers discussing general theoretical or technical issues [10].
The primary objective of this survey is to provide a comprehensive analysis of where the many applications of PSA have been, and where they are going. To ensure a manageable survey scope aligned with a precise, foundational definition of business analytics [12], we focus on a prominent subset of PSA that we call Data-Driven PSA (DPSA, pronounced ”deepsa”). As our analysis in Section 4 shows, DPSA uniquely encompasses the fully realized data-driven, automated decision workflows envisioned by founding figures within business analytics. The pursuit of our primary objective of providing a comprehensive survey is decomposed into the following three research questions:
  1. 1.
@@
    We derive 10 application domains and subsumed problem types and analyse their temporal development, forming a comprehensive, updated overview of current DPSA use cases.
  • 2.
@@
    We derive 5 method types along with a taxonomy of subsumed methods and analyse their temporal development, forming a comprehensive, updated overview of methods found within existing DPSA applications.
  • 3.
@@
    We derive 2 generic workflow patterns and subsumed concrete workflow patterns, and provide a statistical overview of different method type usage combinations, including their temporal development, forming a comprehensive, updated overview of how DPSA application methodologies are structured.
  • 4.
@@
    Finally, we derive 4 possible research directions reflecting the current state of the art, based on an analysis of frequently recurring issues and temporal trends within DPSA applications.
@@
The remainder of this paper is structured as follows: First, we provide a systematic analysis of how the present survey marks a relevant scientific contribution compared to existing PSA surveys in Section 2. Section 3 covers relevant background for readers unfamiliar with PSA. Then Section 4 provides a definition of DPSA, to delimit the scope of this survey. Sections 5-7 move on to present the chosen survey methodology, choices made as part of the survey process, along with a statistical overview of included papers, respectively. Section 8, Sections 9-11, and Section 12 address each research question in turn, providing a comprehensive framework of problem and solution types found within existing application research, along with a number of possible future research directions. We provide concluding remarks in Section 13.
2 Comparison to existing PSA surveys
@@
Several surveys about PSA have been published within recent years. An analysis of feature overlaps between existing surveys and this one can be found in Table 1, with papers sorted by publication year from left to right.
Feature / Paper [13] [10] [9] [14] [15] [16] [17] [18] [19] [20] [21] [22] [23] This survey
Papers from 2020s included
Applications the primary focus
Covers more than one sector
Analysis of application domains (✓) (✓) (✓)
Analysis of problem types (✓) (✓) (✓)
Analysis of utilized methods (✓)
Analysis of workflow patterns
Analysis of temporal trends
@@
Table 1: This survey’s overlaps with existing PSA surveys, with papers sorted by publication year from left to right. (✓) denotes significantly limited fulfilment. [24] and [25] are respectively older versions of [23] and [9], and are therefore left out for the sake of brevity.
@@
First, we note that a number of existing surveys don’t cover any papers from this decade [13, 10, 9, 14]. As our analysis in Section 7 shows, 76.9% of our surveyed papers were published within the 2020-2024 year range, with the number of papers having accelerated significantly over the past three years. Looking solely at the recency of content, this survey therefore provides substantially updated knowledge.
@@
Furthermore, only a few past surveys have had a primary focus on PSA applications, with others focusing on systems and tools [10, 22], methodologies [9, 17], or sociotechnical system analysis [23, 13]. Multiple surveys indeed focus primarily on applications, but they either limit their scope to studying specific application domains or analysis methods utilized therein [15, 16, 18, 17, 21, 20].
@@
Remaining application-oriented papers limit their scope to a significantly limited subset of applications by more or less explicit survey-methodological choices [14, 19]. One of these papers survey ”various” PSA applications, as its title aptly conveys, yet only 21 in total, and without specifying how these papers were found or selected [14]. The other paper explicitly limits its search query to only include PSA papers intersecting with keywords under the topic of ”sustainable operations research” [19]. These methodological choices, excluding the larger portion of existing PSA papers, severely limit how comprehensive analyses of PSA application domains and problem types can be, which is reflected by parenthesized check marks in Table 1.
@@
The most recent PSA survey paper [23] only covers the ”more prevalent” (in terms of the number of papers) application domains and problem types, or ”affordances” in the chosen nomenclature, which is also reflected in Table 1. The aforementioned survey also mentions a number of overall PSA method categories and example methods. These are however not derived from its own set of survey papers, but from Lepenioti et al.’s survey from 2020 [9].
@@
In addition to being comprehensive unlike previous endeavours, this survey marks a relevant research contribution by introducing novel analysis dimensions, namely an analysis and taxonomy of how solution methods are combined into PSA workflows. Although others have considered the relevance of studying this topic [23], only individual workflow patterns have been discussed fragmentarily within PSA research so far (e.g., [26]). Furthermore, while existing surveys sometimes include statistics on the number of papers published per year (e.g., [9]), none provide deeper temporal insights into the PSA application research field’s development over time in terms of encountered problem types or solution methods - another gap closed by the present survey.
3 Background on PSA
@@
As described in in Davenport and Harris’ seminal book on the subject, Competing on Analytics: The New Science of Winning [12], problem solving being evidence-based more concretely entails that business decisions should be driven by the ”extensive use of data, statistical and quantitative analysis, explanatory and predictive models, and fact-based management”, enabled by a set of technologies and processes [12,  p. 7].
@@
4 PSA definitions and survey scope
To ensure a well-defined scope, it was ultimately deemed necessary to use a more specific version of Definition 2 in this survey. Throughout history, many prescriptive approaches have been used within human organizational endeavours: from ancient astronomy using the constellations to time when to harvest the crops [34,  p. 4], to modern operations research arriving to support military efforts during the Second World War [33]. Even disregarding definitional issues, a single survey paper cannot realistically aspire to cover everything written within the scope of Definition 2 since the dawn of history, let alone entire fields like operations research, which came to be more than 80 years ago [35], with a corresponding number of papers to boot.
Instead, we have chosen to refine our scope to focus on prescriptive approaches following a data-driven, automated paradigm in terms of solution methods. With this focus, we answer recent calls for more data-driven PSA research [11, 9, 36]. As we will argue next, the aforementioned PSA paradigm is furthermore characteristic of the business analytics and technological context in which the ”PSA” term originally arose during the past decade (cf. Section 3).
To start, consider the defining characteristics of analytics given in Davenport and Harris’ seminal book [12]. First, analytics is literally understood as a subset of business intelligence: ”a set of technologies and processes that use data to understand and analyse business performance” (our emphasis) [12,  p. 8]. Analytics is thus defined as a fundamentally data-driven endeavour. Second, while it is recognized in the same book that analytics may technically either be input to human decisions, and thus subjected to manual intervention, or be fully automated, it is also stated that any ”rational individual” (our paraphrase) using analytics in the current age would prefer to rely on information technology over pen and paper [12,  p. 8], signifying a preference for automation when possible. Apart from this singular perspective on the matter, many existing definitions of and rationales behind business analytics explicitly mention some combination of being data-driven and leveraging automation as part of its key traits, as summarized in [3].
In addition, note how the ”PSA” term arose within the technological context of the 2010s (cf. Section 3): a time period during which increasingly scalable computing systems for (Big) data-driven, automated analytics had a momentous impact on industry and society [37, 38], and a time period during which the data-driven analytics ecosystem underwent rapid expansion in terms of user base and available tools [39], offering increasing levels of workflow automation [40].
@@ 4 PSA definitions and survey scope
Data-driven prediction, exemplified by data mining and ML methods, accords with original characterizations of business analytics discussed in this section, while automatic prescription, exemplified by mathematical optimization, accords with how the value proposition of PSA was originally formulated by IBM (cf. Section 3): finding the best decision options within problem spaces otherwise too large or complex to search manually.
@@
5 Survey methodology
This section outlines the utilized survey methodology, taking its onset in the recommendations followed by previous surveys within PSA [23, 9]. Sections 6 and 7 provides a statistical summary of the search process and identified papers, respectively. Throughout, we follow the PRISMA 2020 guidelines for systematic literature review reporting [41].
5.1 Databases
@@ 5.2 Search queries
5.3 Selection criteria
To narrow down initial search results to a subset adequately addressing the research questions (cf. Section 1), a set of selection criteria was developed, used for screening the initial set of records in several phases. Within the context of this survey, we will refer to the remaining set of records after screening with these criteria as the set of surveyed papers, or just papers for short when unambiguous. The set of selection criteria is as follows, to be explained in the remainder of this section:
  1. 1.
@@ 5.3 Selection criteria
As for C5, the requirement of developing an end-to-end application means that the paper must analyse a concrete problem solved with concrete methods. For instance, it is insufficient for a paper to simply discuss hypothetical applications of PSA within some application domain on a purely conceptual level, or to propose ”machine learning” as a generic solution to a specific real-world problem without further details. This criterion is introduced in order to be able to assess adherence to Definition 3 and adequately address the research questions (cf. Section 1).
@@
6 Survey process
This section describes the process of how the survey procedure was executed, including choices made during screening. Using the databases, queries and selection criteria of Sections 5.1, 5.2, and 5.3, the screening process for obtaining the final set of surveyed papers proceeded as depicted in Figure 2. This figure depicts a flow diagram according to the PRISMA 2020 guidelines for systematic reviews [41], summarizing the number of records and choices made during each screening phase.
Figure 2: Flow diagram of the executed survey procedure, using the PRISMA 2020 format [41]. Each box in the left chain represents a phase in a sequential screening workflow, along with its number of input records. The boxes to the right summarize how many records were excluded in each screening phase, as well as the reasons why, with respect to the selection criteria of Section 5.3.
@@ 6 Survey process
A few notes on choices made during each phase are in order, and the reader is advised to confer Figure 2 as needed. For the outlet screening phase, records excluded due to C2 were PhD theses, workshop papers, and papers from the Computing Research Repository (CoRR) which were not peer-reviewed. Exclusions due to C3 and C4 followed the objective metrics described in Section 5.3 by looking up the relevant information in the chosen online registers, with the larger part of records being excluded due to the outlet being low-impact (C4).
@@
7 Survey statistics
A total of 104 papers were obtained from executing the survey procedure. Figure 3 summarizes the distribution of papers per year, which spans the entire 2015-2024 range. There is evidently a pronounced skew towards more recent years, which might be attributed to the increasing popularity and impact of DPSA applications within PSA research, with the 2020-2024 range making up about 76.9% of all papers. Supplementary survey statistics can be found in A.